OpenStudy (anonymous):

how to find the lateral area of a right rectangular prism with width 5 length 6 and height 9

OpenStudy (anonymous):

I got 258 but I think that was surface area..

OpenStudy (anonymous):

2(30+54+45) 2(129) LA=258 I think it's surface area though.. not sure how to find LA

OpenStudy (hoblos):

lateral area of a right rectangular prism = perimeter * height = (2*5 + 2*6) * 9 = 22*9 = 198

OpenStudy (anonymous):

how did you know which numbers were the perimeter?

OpenStudy (hoblos):

it is the perimeter of the base of width and 5 length 6

OpenStudy (phi):

The lateral area is the area of the sides. It is the surface area minus the area of the bases (top and bottom faces). It is arbitrary which face you call the "base" , so the only way to know is to interpret the base as the width by length. (The height goes up, so it is not part of the base). But if someone tipped it over, and told you the base was 9 x 5, the lateral area would change.